OpenStudy (anonymous):

write this expression 32^-2/5 in radical notation and then evaluate expression

OpenStudy (anonymous):

so I think that it would be 1/32^2/5

OpenStudy (anonymous):

\[1/32^{2/5}\]

OpenStudy (anonymous):

then i am lost after that

OpenStudy (anonymous):

then I think that yo have to simplify with fractional power so that means that it would be\[1/\sqrt[5]{32}\]

OpenStudy (anonymous):

so teh n2*2*2*2*2=32 so woudl that mean that it would be 1/2 for the answer

OpenStudy (anonymous):

I'd think that the question requires squaring the denominator before finding it's fifth root, maybe: 1/(1024^(1/5))

OpenStudy (anonymous):

but would you simplify the denominator of 32 so that means that it would be \[1/\sqrt[5]{2^{5}}\]

OpenStudy (anonymous):

Yes, if you can cancel out the power and root to pull out two, giving you 1/2, you'd have an answer